daisyy:

Which monomials are perfect cubes? Check all that apply. 3x3 8x12 27x9 100x3 125x5

jhonyy9:

do you know what mean monomials ?

daisyy:

yessirski

jhonyy9:

ok so in case of what choices do you see (ax)^3

jhonyy9:

ok step by step 3x^3 so from here 3 is cube ?

daisyy:

im so sorry but i dont pay attention in class but im starting to now its just hard rn lol um help

jhonyy9:

no the second one choice 8x^12 8 = 2^3 - cube x^12 = (x^4)^3 - cube in this way

jhonyy9:

from this result 8x^12 is perfect cubes

daisyy:

im rlly tryna understand but its hard to focus

jhonyy9:

@Laylalyssa any idea ?

Laylalyssa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 @Laylalyssa any idea ? \(\color{#0cbb34}{\text{End of Quote}}\) nah

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Laylalyssa \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 @Laylalyssa any idea ? \(\color{#0cbb34}{\text{End of Quote}}\) nah \(\color{#0cbb34}{\text{End of Quote}}\) to explain it more understandably

daisyy:

its fine i got it